The Joy of Conics

or How to Investigate and Classify

Ellipses, Parabolas, and Hyperbolas

In Memory of Euclid

Apollonius from Pergae^{1} and Archimedes from Syracuse

Abstract. The various types of conic sections are introduced including a justification for their respective names ellipsis, parabola, hyperbola. Several examples are given.

2010 Mathematics Subject Classification: Primary 11E16; Secondary 11D09, 11E04, 15A63. Keywords and Phrases: Conic section, ellipsis, parabola, hyperbola.

This is just a test for page 1.

L^{A}TEX 2ϵ

1Supported by the Science Exchange Programme of the Universities of Alexandria and Pergamon

Documenta Mathematica 23 (2018) 1–5

2 Apollonius and Archimedes

This is just a test for page 2. L^{A}TEX 2ϵ

Documenta Mathematica 23 (2018) 1–5

The Joy of Conics 3 This is just a test for page 3.

L^{A}TEX 2ϵ

Documenta Mathematica 23 (2018) 1–5

4 Apollonius and Archimedes

This is just a test for page 4. L^{A}TEX 2ϵ

Documenta Mathematica 23 (2018) 1–5

This is just a test for page 5.

L^{A}TEX 2ϵ

Apollonius from Pergae Department of Mathematics University of Pergamon Pergamon Western Asia Minor, Persia diana@math.uni-pergae.pe Archimedes from Syracuse King Hieron’s Palace Heureka Avenue Syracuse Sicily, Greece urania@math.king-syr.sic

Documenta Mathematica 23 (2018) 1–5

6

Documenta Mathematica 23 (2018)